# images/svg/

### v1

@@ Abstract
Inspection of complex underwater structures with tethered underwater vehicles is often hindered by the risk of tether entanglement. We propose REACT (real-time entanglement-aware coverage path planning for tethered underwater vehicles), a framework designed to overcome this limitation. REACT comprises a fast geometry-based tether model using the signed distance field (SDF) map for accurate, real-time simulation of taut tether configurations around arbitrary structures in 3D. This model enables an efficient online replanning strategy by enforcing a maximum tether length constraint, thereby actively preventing entanglement. By integrating REACT into a coverage path planning framework, we achieve safe and optimal inspection paths, previously challenging due to tether constraints. The complete REACT framework’s efficacy is validated in a pipe inspection scenario, demonstrating safe, entanglement-free navigation and full-coverage inspection. Simulation results show that REACT achieves complete coverage while maintaining tether constraints and completing the total mission $20\%$ faster than conventional planners, despite a longer inspection time due to proactive avoidance of entanglement that eliminates extensive post-mission disentanglement. Real-world experiments confirm these benefits, where REACT completes the full mission, while the baseline planner fails due to physical tether entanglement.
@@ I Introduction
Numerous coverage path planner (CPP) algorithms are proposed in the literature for inspection-related tasks with untethered systems [4, 5, 6]. In essence, they compute distance-optimal paths for a thorough inspection of 3D structures [7, 8, 9]. Additionally, exploration path planners are employed to determine the next viewpoints for mapping unknown terrains [10, 11]. However, these path-planning methods are restricted to untethered systems, as they do not account for possible entanglements with the surroundings. Hence, the literature still lacks path-planning algorithms that directly address tether-related challenges.
Figure 1: Overview of the REACT inspection framework. Offline, a SDF map is generated from a point cloud, and an off-the-shelf CPP [8] is used to compute an optimal waypoint sequence. During operation, a tether model ($\mathbf{P}_{\text{tether}}$) is used to ensure tether length is not exceeded, while a MPC applies the optimal wrench $\mathbf{u}$ to the ROV.
@@
  • A computationally efficient tether model that computes the tether configuration using the SDF map of arbitrary underwater structures.
@@
  • Integration with an off-the-shelf CPP along with MPC, rendering optimal inspection trajectory.
  • Demonstration of the framework in simulation and real-world tests, showcasing its ability to ensure safe and time-efficient inspection.
@@ II State of the art
Initial efforts to develop entanglement-aware path planners were often focused on two-dimensional environments and employed offline computation strategies. For instance, in [13] and [14], methods were proposed for planning paths in 2D to cover a predefined set of waypoints while considering tether constraints. Notably, in [14], the path planning problem for a ROV was formulated as a mixed-integer programming problem. In that approach, a traveling salesman problem (TSP) was first solved to find an optimal waypoint sequence, and then homotopic constraints were incorporated during path generation to minimize the likelihood of tether entanglement. While effective for predefined scenarios, these offline methods lack the adaptability required for dynamic environments. A more recent approach to the 2D coverage path planning problem for tethered robots was presented in [15], where spanning tree-based offline optimization was employed.
To address the need for real-time adaptability, subsequent research explored online path planning algorithms, still primarily in 2D. In [16], a homotopy-augmented topological approach combined with graph search techniques was introduced, allowing for dynamic adjustments to the path based on environmental perception. Similarly, in [17], a hybrid A* variant utilizing a modified tangent graph was developed. This method efficiently plans curvature-constrained paths for tethered robots subject to winding angle constraints, demonstrating guarantees and providing simulation results for online entanglement avoidance.
The complexity of tether management increases significantly when coordinating multiple robots. Foundational work in this area was laid in [18], followed by methods proposed in [19]. More recently, in [20], an efficient online path planner for multi-robot systems was presented. In this method, a homotopy-based high-level planner was integrated with trajectory optimization and smoothing techniques to generate entanglement-free paths. However, despite its online capability, this approach remains constrained to 2D environments.
The online planners described above represent a significant step towards real-time tether management; however, they are limited to 2D environments. Moreover, while preventive paths to avoid entanglement can be planned, strategies for path planning once tether entanglement has already occurred are not provided.
@@
The increased complexity of 3D multi-robot scenarios was addressed in [24], where earlier 2D work was extended to three dimensions. Further advancements were introduced in [25] and [26], where path planning strategies explicitly considering the topological constraints imposed by multiple interacting tethers in 3D were proposed. While these methods advance the state of the art in multi-robot coordination, they are generally designed for offline computation and are not suited for online path planning where real-time implementation is essential.
In summary, existing path planners that account for tether constraints often face limitations for practical online CPP in complex 3D settings. Many are too computationally intensive for real-time use [14, 24, 25, 26], lack integrated tether-aware CPP frameworks, or rely on simplifying assumptions such as 2D environments or basic obstacle shapes [16, 17, 20], hindering generalization to real-world inspection tasks.
To address these limitations, a novel approach named REACT is proposed, which enables real-time, entanglement-aware path planning in arbitrary 3D environments.
@@ III Overall REACT Framework
REACT framework, as depicted in Fig. 1, consists of two main components: an offline planning phase and an online execution phase. In the offline phase, the environment is provided as a point cloud, including the object to be inspected. Then, a SDF map is generated from this point cloud using the nvblox library [27]. Subsequently, the extracted point cloud of the structure under inspection is processed via FC-Planner [8] to compute an optimal waypoint sequence, generating a path that renders full inspection coverage while disregarding the tether constraints.
In the online phase, tether constraints are handled by an entanglement-aware replanner that ensures that the maximum tether length is not exceeded due to entanglement. This is achieved using our developed tether model, which continuously updates based on the current state of the tether and the new position of the ROV, denoted as $\textbf{p}_{\mathrm{rov}}$. The online planner then provides the reference state to an MPC controller, which computes and applies the optimal wrench to the ROV.
@@ IV-A Tether model description
Let the tether path at time $t$ be denoted by $\mathbf{P}_{\mathrm{tether}}(t)=\{\mathbf{p}_{i}(t)\}_{i=1}^{n}$, where each node $\mathbf{p}_{i}(t)\in\mathbb{R}^{3}$ represents the position of the $i$-th node in 3D space at time $t$, and $n$ is the total number of nodes in the tether path. The ROV position at time $t$, denoted by $\mathbf{p}_{\mathrm{rov}}(t)\in\mathbb{R}^{3}$, is appended at the end $\mathbf{p}_{n+1}(t)$, ensuring that the ROV’s position is included in the tether path.
The proposed tether model then iteratively computes the equivalent taut-tether path through sequential shortcutting operation. Fig. 2 illustrates an example of the shortcut operation. Starting at the last node ($\mathbf{p}_{j}(t)=\mathbf{p}_{n}(t)$), the algorithm attempts to shortcut the path segment between each pair of nodes $(\mathbf{p}_{i}(t),\mathbf{p}_{j}(t))$ where $i>j$. If the line of sight between $\mathbf{p}_{i}(t)$ and $\mathbf{p}_{j}(t)$ is collision-free, as determined via an SDF map ($\mathcal{M}_{sdf}$), the intermediate nodes are replaced with a straight segment sampled at known resolution $\delta$. Conversely, if the line of sight encounters a collision, $j$ shifts to the preceding node, and the collision-free line of sight is rechecked. This process is repeated until a collision-free line of sight is found, and the intermediate nodes are then replaced.
Finally, after the shortcutting step, a pulling operation is applied to each node, moving it incrementally toward the tether endpoint $\mathbf{p}_{n+1}(t)$. This ensures that nodes are not left stuck in the cavities of non-convex obstacles. The shortcutting and pulling operations are applied iteratively until convergence, resulting in a taut and collision-free tether path $\mathbf{P}_{\mathrm{tether}}(t+1)$. The full procedure is described in Algorithm 1.
@@
appendPath($\mathbf{P}_{\mathrm{tether}}(t)$, $\mathbf{p}_{\mathrm{rov}}(t)$);
for $i\leftarrow\mathrm{len}(\mathbf{P}_{\mathrm{tether}}(t))-1$ to 00 do
       for $j\leftarrow i-1$ to 00 do
@@
                         pullNode($\mathbf{P}_{\mathrm{tether}}(t)[j]$, $\mathbf{P}_{\mathrm{tether}}(t).end()$, $\delta$);
@@ V Entanglement-Aware Path Planner
At each time step $t$, the planner receives the current estimated tether path, $\mathbf{P}_{\mathrm{tether}}(t)$, from the tether model. It also maintains an ordered list of target waypoints, $\mathbf{W}=\{\mathbf{p}_{\text{waypoint}}(k)\}_{k=1}^{m}$, where each $\mathbf{p}_{\text{waypoint}}(k)$ specifies a 3D position the ROV must reach sequentially. The current position of the ROV, $\mathbf{p}_{\text{rov}}(t)$, and the current waypoint index, $k$, are also maintained as part of the planner’s state.
The current tether length $L_{\text{tether}}(t)$ is computed from the tether path and compared against the maximum allowable tether length $L_{\max}$ to determine whether re-planning is necessary.
@@
This dual-mode approach ensures that the vehicle maintains a safe tether configuration while effectively progressing through its mission waypoints. By continuously switching between nominal waypoint tracking and recovery behaviors in response to tether length measurements, the planner enables real-time, entanglement-aware path planning. The overall entanglement-aware planning framework is summarized in Algorithm 2.
@@
$L_{\text{tether}}(t)\leftarrow\text{computeLength}\bigl{(}\mathbf{P}_{\mathrm{tether}}(t)\bigr{)}$;
if $mode=\mathrm{NORMAL}$ then
@@
             $\mathbf{P}_{\text{recovery}}\leftarrow\text{deEntanglementSearch}\bigl{(}\mathbf{P}_{\mathrm{tether}}(t),\mathbf{W}[k]\bigr{)}$;
@@
return $\mathbf{p}_{\mathrm{target}}$;
@@ V-A Disentanglement path search
To implement the disentanglement behavior described above, the recovery path search algorithm (Algorithm 3) is employed. This algorithm takes as input the current tether path $\mathbf{P}_{\mathrm{tether}}(t)$, the current target waypoint $\mathbf{W}[k]$, and the maximum allowable tether length $L_{\max}$. Its goal is to generate a recovery path $\mathbf{P}_{\mathrm{recovery}}$ that guides the ROV safely toward the waypoint while ensuring that the tether length constraint is respected.
@@
For each pivot, two path segments are constructed. The first segment, $\mathbf{P}_{\mathrm{r}1}$, corresponds to the reversed portion of the tether from the pivot $\mathbf{p}_{\mathrm{pivot}}$ to the ROV’s current position, representing the retraced section of the tether. The second segment, $\mathbf{P}_{\mathrm{s}(n-i)}$, is a newly planned shortest path connecting the pivot to the target waypoint $\mathbf{W}[k]$, which can be computed using sampling-based methods such as RRT*. These segments are then concatenated with the tether path from $\mathbf{p}_{0}$ up to the pivot, forming an augmented candidate path $\mathbf{P}_{\mathrm{aug}}$.
This candidate path $\mathbf{P}_{\mathrm{aug}}$ is used to simulate the tether configuration and estimate its length $L_{\mathrm{tether}}$ at the next time step. The total tether length $L_{\mathrm{tether}}$ is then compared against the maximum allowable length $L_{\max}$. If the length constraint is satisfied, the recovery path $\mathbf{P}_{\mathrm{recovery}}=\mathbf{P}_{\mathrm{r}1}\cup\mathbf{P}_{\mathrm{s}(n-i)}$ is considered feasible and returned.
If no feasible path is found after examining all possible pivot points, the algorithm defaults to a fallback strategy where the tether length constraint is treated as a soft limit. In this case, the recovery path consists solely of the shortest path from the ROV’s current position directly to the target waypoint, temporarily disregarding tether length constraints.
The resulting recovery path provides a safe and viable trajectory for the ROV to follow during disentanglement operations by combining the reversed tether segment (if applicable) with the newly planned path toward the waypoint, as illustrated in Figure 3.
@@
       $\mathbf{P}_{\text{aug}}\leftarrow\text{concatenate}\bigl{(}\mathbf{P}_{\text{tether}}(t)[0:i],\mathbf{P}_{\text{s}(n-i)}\bigr{)}$;
@@
       if $L_{\text{tether}}\leq L_{\max}$ then
@@
if not $found\_feasible$ then
@@
$\mathbf{P}_{\text{recovery}}\leftarrow\mathbf{P}_{\text{r}1}\cup\mathbf{P}_{\text{r}2}$;
return $\mathbf{P}_{\text{recovery}}$;

### v2

@@ Abstract
Inspection of underwater structures with tethered underwater vehicles is often hindered by the risk of tether entanglement. We propose REACT (real-time entanglement-aware coverage path planning for tethered underwater vehicles), a framework designed to overcome this limitation. REACT comprises a computationally efficient geometry-based tether model using the signed distance field (SDF) map for accurate, real-time simulation of taut tether configurations around arbitrary structures in 3D. This model enables an efficient online replanning strategy by enforcing a maximum tether length constraint, thereby actively preventing entanglement. By integrating REACT into a coverage path planning framework, we achieve safe and entanglement-free inspection paths, previously challenging due to tether constraints. The complete REACT framework’s efficacy is validated in a pipe inspection scenario, demonstrating safe navigation and full-coverage inspection. Simulation results show that REACT achieves complete coverage while maintaining tether constraints and completing the total mission $20\%$ faster than conventional planners, despite a longer inspection time due to proactive avoidance of entanglement that eliminates extensive post-mission disentanglement. Real-world experiments confirm these benefits, where REACT completes the full mission, while the baseline planner fails due to physical tether entanglement.
@@ I Introduction
Numerous coverage path planner (CPP) algorithms are proposed in the literature for inspection-related tasks with untethered systems. In essence, they calculate approximate distance-optimal paths for a thorough inspection of 3D structures [8, 13, 5]. In addition, exploration path planners are employed to determine the next points of view to map unknown terrains [12]. However, these path-planning methods are restricted to untethered systems, as they do not account for possible entanglements with the surroundings. Hence, the literature still lacks tether-aware inspection planners.
Figure 1: Overview of REACT: Offline, an SDF map is generated from a point cloud, and an off-the-shelf CPP [13] is used to compute a near distance-optimal waypoint sequence. Online, REACT provides an entanglement-free next waypoint to the controller.
@@
  • A computationally efficient tether model that computes the entangled tether configuration.
@@
  • A complete entanglement-aware inspection framework integrating CPP, SDF-based mapping, and model predictive control (MPC) to enable safe asset inspections.
  • Demonstration of the framework in simulation and real-world tests, showcasing safe and full coverage inspection.
@@ II State of the art
Initial efforts in entanglement-aware path planning focused on 2D environments and relied on offline computation. Notably, [18] introduced the non-entangling traveling salesman problem, formulating the ROV path planning problem as a mixed-integer program that simultaneously determines the waypoint sequence and the path’s homotopy class, guaranteeing an entanglement-free trajectory. Similarly, [19] addressed 2D waypoint coverage under tether constraints. While effective for predefined scenarios, these offline approaches lack adaptability in dynamic environments. More recently, [22] proposed a spanning tree-based offline optimization for 2D coverage.
To address the need for real-time adaptability, subsequent research explored online path planning algorithms, still primarily in 2D. In [16], a homotopy-augmented topological approach combined with graph search techniques was introduced, allowing for dynamic adjustments to the path based on environmental perception. Similarly, in [11], a hybrid A* variant utilizing a modified tangent graph was developed. This method efficiently plans curvature-constrained paths for tethered robots subject to winding angle constraints, demonstrating guarantees and providing simulation results for online entanglement avoidance. Online cell-based decomposition strategies are also proposed to dynamically handle tether constraints and environmental changes in 2D [26].
The complexity of tether-aware path planning further increases when coordinating multiple robots. Early work in this area is presented in [14], followed by the method proposed in [28]. More recently, in [10], an efficient online path planner for multi-robot systems was presented. In this method, a homotopy-based high-level planner was integrated with trajectory optimization and smoothing techniques to generate entanglement-free paths. Similarly, [27] proposed a solution to the tethered robot pair motion planning problem for tethered robots in 2D, employing a reduced visibility graph to account for tether interactions with polygonal obstacles. However, despite these advances, these approaches remain constrained to 2D environments. Moreover, while preventive paths to avoid entanglement can be planned, strategies for path planning once tether entanglement has already occurred are not provided.
@@
The increased complexity of 3D multi-robot scenarios was addressed in [15], where previous 2D work was extended to three dimensions. Further advances were introduced in [21] and [9], where path planning strategies explicitly considered the topological constraints imposed by multiple interacting tethers in 3D were proposed. While these methods advance the state of the art in multi-robot coordination, they are generally designed for offline computation and are not suited for online planning where real-time performance is needed.
In summary, existing path planners that account for tether constraints often face limitations for practical online CPP in complex 3D settings. Many are too computationally intensive for real-time use [19, 15, 21, 9], lack integrated tether-aware CPP frameworks, or rely on simplifying assumptions such as 2D environments or basic obstacle shapes [16, 11, 10], hindering generalization to real-world inspection tasks. To address these limitations, REACT is proposed, which enables real-time, entanglement-aware path planning in arbitrary 3D environments.
@@ III Overall REACT Framework
REACT framework, as depicted in Fig. 1, consists of two main components: an offline planning phase and an online execution phase. In the offline phase, the environment is provided as a point cloud, including the object to be inspected. Then, a SDF map is generated from this point cloud using the nvblox library [20]. Subsequently, the extracted point cloud of the structure under inspection is processed via FC-Planner [13] to compute an optimal waypoint sequence (referred to as nominal waypoints), generating a path that renders full inspection coverage while disregarding the tether constraints.
In the online phase, tether constraints are managed by an entanglement-aware replanner that ensures that the maximum tether length is not exceeded. The proposed REACT framework incorporates a modular online layer that can be seamlessly integrated with any existing offline coverage or path planner. By decoupling coverage planning from tether-constrained replanning, the framework transforms an otherwise highly complex and computationally intractable joint problem into two tractable subproblems: offline coverage planning and fast online tether management. This separation not only enables the use of standard off-the-shelf coverage planners, but also greatly simplifies integration with other planning frameworks.
Operating online provides further advantages: it allows the system to handle operational disturbances and model uncertainties that cannot be fully anticipated during offline planning. Moreover, it naturally supports scenarios where the map expands during operation, such as in exploration or when new obstacles are detected, enabling adaptive planning in previously unknown environments. The online management-aware planner then provides entanglement-free reference trajectories that are tracked by a MPC controller, which translates them into optimal wrench commands for the ROV. This ensures that the vehicle follows the re-planned paths while respecting dynamics and actuation limits.
@@ IV-A Tether model description
To efficiently compute the tether configuration around obstacles, the continuous tether is discretized into a finite set of nodes, separated by a fixed resolution $\delta$. Each node represents a point along the tether and collectively approximates the overall 3D shape of the tether. This discretization enables computationally efficient collision checking and shortcutting.
Let the tether path at time $t$ be denoted by $\mathbf{P}_{\mathrm{tether}}(t)=\{\mathbf{p}_{i}(t)\}_{i=1}^{n}$, where each node $\mathbf{p}_{i}(t)\in\mathbb{R}^{3}$ represents the position of the $i$-th node in 3D space at time $t$, and $n$ is the total number of nodes in the tether path. The ROV position at time $t$, denoted by $\mathbf{p}_{\mathrm{rov}}(t)\in\mathbb{R}^{3}$, is appended at the end $\mathbf{p}_{n+1}(t)$. The proposed tether model then iteratively computes the equivalent taut-tether path through sequential shortcutting operation. Fig. 2 illustrates an example of the shortcut operation. Starting at the last node ($\mathbf{p}_{j}(t)=\mathbf{p}_{n}(t)$), the algorithm attempts to shortcut the path segment between each pair of nodes $(\mathbf{p}_{i}(t),\mathbf{p}_{j}(t))$ where $i>j$. If the line of sight between $\mathbf{p}_{i}(t)$ and $\mathbf{p}_{j}(t)$ is collision-free, as determined via an SDF map ($\mathcal{M}_{sdf}$), the intermediate nodes are replaced with a straight segment sampled at known resolution $\delta$. Conversely, if the line of sight encounters a collision, $j$ shifts to the preceding node, and the collision-free line of sight is rechecked. This process is repeated until a collision-free line of sight is found, and the intermediate nodes are then replaced. Finally, after the shortcutting step, a pulling operation is applied to each node, moving it incrementally toward the tether endpoint $\mathbf{p}_{n+1}(t)$. This ensures that nodes are not left stuck in the cavities of non-convex obstacles. The shortcutting and pulling operations are applied iteratively until convergence, resulting in a taut and collision-free tether path $\mathbf{P}_{\mathrm{tether}}(t+1)$. The full procedure is described in 1.
Remark 1
Sparse environment reconstructions may create gaps in the SDF, causing obstacles to appear partially free. This can lead to unrealistic tether shortcutting through obstacles unless sufficient voxel resolution is applied.
Remark 2
The tether discretization resolution $\delta$ should be chosen consistently with the SDF voxel resolution to ensure reliable collision detection.
@@
1ex
[…]
for $i\leftarrow\mathrm{len}(\mathbf{P}_{\mathrm{tether}}(t))-1$ to 0 do
 for $j\leftarrow i-1$ to 0 do
@@
[…]
@@ V Entanglement-Aware Path Planner
At each time step $t$, the planner receives the current estimated tether path, $\mathbf{P}_{\mathrm{tether}}(t)$, from the tether model. It also maintains an ordered list of target waypoints, $\mathbf{W}=\{\mathbf{p}_{\text{waypoint}}(k)\}_{k=1}^{m}$, where each $\mathbf{p}_{\text{waypoint}}(k)$ specifies a 3D position the ROV must reach sequentially. The current position of the ROV, $\mathbf{p}_{\text{rov}}(t)$, and the current waypoint index, $k$, are also maintained as part of the planner’s state. The current tether length $L_{\text{tether}}(t)$ is computed from the tether path and compared against the maximum allowable tether length $L_{\max}$ to determine whether replanning is needed.
@@
This dual-mode approach ensures that the vehicle maintains a safe tether configuration while effectively progressing through its mission waypoints. By continuously switching between nominal waypoint tracking and recovery behaviors in response to tether length measurements, the planner enables real-time, entanglement-aware path planning. The overall planning framework is summarized in Algorithm 2.
@@
[…]
1exif $mode=\mathrm{NORMAL}$ then
@@
    $\mathbf{P}_{\text{recovery}}\leftarrow\text{disEntanglementSearch}\bigl(\mathbf{P}_{\mathrm{tether}}(t),\mathbf{W}[k]\bigr)$;
@@
1exreturn $\mathbf{p}_{\mathrm{target}}$;
@@ V-A Disentanglement path search
To implement the disentanglement behavior described above, the recovery path search algorithm (Algorithm 3) is employed. This algorithm takes as input the current tether path $\mathbf{P}_{\mathrm{tether}}(t)$, the current target waypoint $\mathbf{W}[k]$, and the maximum allowable tether length $L_{\max}$. Its goal is to generate a recovery path $\mathbf{P}_{\mathrm{recovery}}$ that guides the ROV safely toward the waypoint while respecting the tether length constraint.
@@
For each pivot, two path segments are constructed. The first segment, $\mathbf{P}_{\mathrm{r}1}$, corresponds to the reversed portion of the tether from the pivot $\mathbf{p}_{\mathrm{pivot}}$ to the ROV’s current position, representing the retraced section of the tether. The second segment, $\mathbf{P}_{\mathrm{s}(n-i)}$, is a newly planned shortest path connecting the pivot to the target waypoint $\mathbf{W}[k]$, which can be computed using sampling-based methods such as RRT*. These segments are then concatenated with the tether path from $\mathbf{p}_{0}$ up to the pivot, forming an augmented candidate path $\mathbf{P}_{\mathrm{aug}}$, which is used to simulate the tether configuration and estimate its length $L_{\mathrm{tether}}$ at the next time step. The total tether length $L_{\mathrm{tether}}$ is then compared against the maximum allowable length $L_{\max}$. If the length constraint is satisfied, the recovery path $\mathbf{P}_{\mathrm{recovery}}=\mathbf{P}_{\mathrm{r}1}\cup\mathbf{P}_{\mathrm{s}(n-i)}$ is considered feasible.
If no feasible path is found after examining all possible pivot points, the algorithm defaults to a fallback strategy where the tether length constraint is treated as a soft limit. In this case, the recovery path consists solely of the shortest path from the ROV’s current position directly to the target waypoint, temporarily disregarding tether length constraints. The resulting recovery path provides a safe and viable trajectory for the ROV to follow during disentanglement operations by combining the reversed tether segment with the newly planned path to the waypoint, as shown in Fig. 3.
@@
[…]
@@
  1exif $L_{\text{tether}}\leq L_{\max}$ then
@@
1exif not found_feasiblefound\_feasible then
@@
[…]
return $\mathbf{P}_{\mathrm{recovery}}$;